ISCE to StaMPS Manual

Version 1.0.1

Author: David Bekaert

Jet Propulsion Laboratory, California Institute of Technology

United States Government Sponsorship acknowledged 20 April, 2018

Contents

1 DISCLAIMER  1
2 Installation  2
3 Introduction  3
4 Setting-up StaMPS processing directories  4
4.1 PS Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5
4.2 SB Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5
5 Additional scripts  6

Chapter 1

DISCLAIMER

This manual and the included scripts call routines of the JPL/Caltech ISCE software to prepare outputs of the ISCE stack processort for ingestion in the open source StaMPS package. It is assumed that the user has already generated the coregisterd stack using the ”SLC” workflow of the stack processors included ISCE contrib folder. At this time, the ISCE to StaMPS package supports only outputs from the topsStack processor.

In additon to citing ISCE and the stack processor, cite the ISCE to StaMPS as:

D. P. S. Bekaert et al. Spaceborne Synthetic Aperture Radar Survey of Subsidence in Hampton Roads, Virginia (USA), Scientific Reports (2017). DOI: 10.1038/s41598-017-15309-5

The provided scripts as part of this ISCE to StaMPS conversion package, are included into the contrib of ISCE and are not supported as part of official ISCE applications. The provided scripts only set-up the assumed StaMPS processing structure. Users should refer back to StaMPS for actual time-series processing. The scripts are provided to you ”as is” with no warranties of corrections. Use at your own risk.

Author: David Bekaert Organization: Jet Propulsion Laboratory, California Institute of Technology United States Government Sponsorship acknowledged.

Chapter 2

Installation

The provided scripts call ISCE and StaMPS functionality, both packages should be installed. In addition the bin folder should be added to your PATH.

Introduction

A set of shell scripts are provided that set-up the StaMPS directory structure using outputs from the ISCE stack processors. Outputs from stripmapApp.py and topsApp.py are not supported as these are pair-by-pair processing apps in ISCE. Only outputs from the topsStack and stripmapStack processors are supported in which the user should have selected the ”SLC” workflow.

The SLC workflow outputs a stack of coregisted SLC’s, baselines grids for each SLC, and geometry files (e.g., los, lon, lat, hgt). Each of these files should a corresponding xml file. In case xml files are needed, use the gdal2isce xml.py script included in ISCE package to generate them.

The following structure is assumed:

   • SLC’s are organized as yyyymmdd/yyyymmdd.slcSLC suffix, where the SLC suffix

is an optional suffix string.

• Baselines  are  organized  as  master slave/master slave  or

     master slave/master slave.txt,  where master and slave are in
     yyyymmdd format.
   • Geom data are stored together as lon.rdrgeom suffix, lat.rdrgeom suffix,
     hgt.rdrgeom suffix and los.rdrgeom suffix, where the geom suffix is an

optional suffix string.

Setting-up StaMPS processing directories

Generate a file called ”input file” at the location where the INSAR master date needs to be generated. The following parameters should be contained in the input file. Required parameters are indicated with (R), optional parameters with (O). The master date can be chosen based on minimizing perpendicular and temporal baseline. The baselines will be automatically re-inverted to the chosen master.

Parameter Name Value Description: R=required,O=optional

source data  slc stack  Do not change (R)
slc stack path                PATH SLC  Path to the SLC stack (R)

slc stack master

yyyymmdd

Master date as to be used in StaMPS processing. Does not need to be the same as the master used for generating the coregistered stack. (R)

slc stack geom path           PATH geom  Path to the geometry directory. (R)

slc stack baseline path       PATH baseline  Path to the baselines directory (R)

range looks

40

Number of range looks multi-looked products (R)

azimuth looks

10

Number of azimuth look multi-looked products (O). azimuth looks over-

rules aspect ratio.
aspect ratio

4

Aspect ratio multi-looked products (R)

lambda                       0.056  Wavelength in meter units (R)
slc suffix                    .full  Suffix string for SLC’s (O)
geom suffix                   .full  Suffix string for geom files (O)

To generate the single master setup type:

make single master stack isce

4.1 PS Processing

If you plan to run the StaMPS PS processing, proceed directly to the ”PS processing” chapter of the StaMPS manual. Unlike the mt prep command detailed in the StaMPS manual use mt prep isce instead.

4.2 SB Processing

If you plan to only run the StaMPS Small Baseline processing, it will save time and disk space to turn the hard-coded flag for the SM interferogram and amplitude file generation to ’n’ in the make single master stack isce code. Even when not generating the interferograms or amplitude files, the required information for the SMALL BASELINES processing will be extracted.

If you did not run the STAMPS PS processing in the INSAR master date directory, then first load baseline info into the matlab workspace. If you already ran the PS processing, you can skip this step:

mt extract info isce
matlab
>>ps load info

To determine which small baseline interferograms to make, in the INSAR master date directory run one of the sb find*.m codes included in StaMPS, e.g.,:

matlab
>>sb find (rho min , Ddiff max , Bdiff max)

Adjust the input parameters according to your data set, where rho min is the minimum coherence based on the maximum temporal Ddiff max and perpendicular Bdiff max decorrelated baselines. There should be no isolated clusters of images. More connections can be made by reducing rho min, or individual connections can be added by editing small baselines.list, which is created by sb find. The connections in small baselines.list can then be plotted with the default StaMPS program:

>>plot sb baselines.

To create the small baseline interferograms listed in small baselines.list in the

INSAR master date directory, run:
make small baselines isce.

Once completed, proceed directly to ”Small Baseline processing” chapter of the StaMPS manual. Unlike the mt prep command detailed in the StaMPS manual use mt prep isce instead.

Chapter 5

Additional scripts

If mdx is intalled, the following command can be used to make snapshots of the interferograms:

get quickview isceisce minrefdem.int